brenda purchased a business for 120,000. and received s building fmv 80,000. land fmv 10,000. and furniture and fixtures fmv 30,000. she spent 10,000. on legal fees to make thr purchase what is the basis for these items
Answer #1

Total cost = 120,000 + 10,000 = 130,000

Total fair market value = 80,000+10,000+30,000 =120,000

.

Building = 130,000*80,000/120,000 = 86,667

Land = 130,000*10,000/120,000 = 10,833

Furniture and fixtures = 130,000*30,000/120,000 = 32,500